OpenStudy (anonymous):

Using complete sentences, explain how solving a literal equation is similar to or different from simplifying an expression such as 6 - 2(52 + 7) ÷ 4.

OpenStudy (anonymous):

First, we will distribute the number in the paranthesis, which means we multiply 52 to 2 and 7 to 2. We will get 104 and 14, but we still put them in paranthesis. This is how our new equation look like: 6-(104+14)÷ 4 Now, we add the number in the ( ). 104 + 14= 118. This is how our new equation look like: 6-(118)÷ 4 We always do the division/ multiplication first. Therefore, we divided 118÷ 4=29.5. And finally, we subtract 6 for 29.5, we get -23.5. This is how the equation look like: 6-29.5=-23.5